OpenStudy (anonymous):

Compute the volume of the solid whose bottom face is the circle x^2+y^2<_1 and every cross section of the solid perpendicular to the x -axis is a square

OpenStudy (anonymous):

@perl

OpenStudy (perl):

it is similar to the last problem we did

OpenStudy (perl):

We are given x^2+y^2=1 Since we want squares perpindicular to x axis we will solve for y in terms of x y= +/- sqrt( 1-x^2) The 'top half' semicircle is the equation y = sqrt(1-x^2) the 'bottom half' semicircle is the equation y = -sqrt(1-x^2) we use the form integral A(x)dx Here A(x) = ((sqrt(1-x^2)-(-sqrt(1-x^2)))^2 which becomes A(x) = (sqrt(1-x^2) +sqrt(1-x^2))^2 A(x) = (2*sqrt(1-x^2))^2 So we have integral A(x)dx on x=-1..1 integral (2*sqrt(1-x^2))^2 dy on x=-1..1